OpenStudy (anonymous):

Add: 45 3 ----- + ------ s^2-49 s-7

OpenStudy (anonymous):

lcm = s^2-49 = 45 + 3(s+7) ---------- s^2-49 = 3s + 66 ------- s^2-49